OpenStudy (anonymous):

Need help Maria invested $450 in a savings account and earned $126 in interest at the end of 7 years. What was the interest rate? A. 3.5% B. 4% C. 4.5% D. 5% 4%

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

I got 4%

OpenStudy (anonymous):

hi chat noir!!I remember you :D

OpenStudy (kittiwitti1):

How're you doing right meow? :3

OpenStudy (anonymous):

Well! ^_^

OpenStudy (kittiwitti1):

That's purrfect~ ♪ ^_^

OpenStudy (anonymous):

You're right It's B

OpenStudy (seratul):

You are right! 126 = 450 × r × 7 126 = 3150r r = 126/3150 r = 0.04 x 100% = 4% R is rate which you need to find.

OpenStudy (anonymous):

Thanks you two!